Environ Systems is a firm that specializes in cleaning environmental damage (waste disposal) and

specialty chemicals. The firm has been traded for only 2 years and gets all of its revenue in the United

States. You have been asked to estimate a cost of equity for the firm and have collected the following

information:

The average regression beta across waste disposal firms is 1.34 and the average market debt to

equity ratio for these firms is 12%; the average regression beta for chemical firms is 1.15 and the

average market debt to equity ratio is 25%. The corporate tax rate is 40%.

Environ Systems’ book value of equity is $ 500 million but the market value of equity is $ 2 billion.

The book (and market) value of debt is $ 500 million.

Estimate the bottom-up levered beta for Environ Systems, assuming that 80% of its value comes

from chemicals and 20% from environmental clean-up. ( I got levered beta 1.436 for waste and 1.3 for chemical ) I think this is wrong.
Answer #1
1) The first step is to unlever the waste disposal firms' and
chemical firms' levered beta using the formula for
unlevering beta.
Unlevered beta = levered beta/(1+ (1-t) (Debt/Equity))
Unlevered beta for waste disposal firms = 1.34/(1+0.60*0.12) = 1.25
Unlevered beta for chemical firms = 1.15/(1+0.60*0.25) = 1.00
2) The next step is to find the unlevered beta appropriate for
Enron. It is the weighted average of the above two betas.
Unlevered beta for Enron = 1.25*20%+1*80% = 1.05
3) Levered beta for Enron = 1.05*(1+0.60*500/2000) = 1.21